OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{\dfrac{P+1}{Q+1}=\dfrac{R}{S}\hspace{.33em}\\~\\ \dfrac{P+2}{Q+2}=\dfrac{1}{2}\hspace{.33em}\\~\\ P+Q=?\hspace{.33em}\\~\\ a.3\hspace{.33em}\\~\\ b.4\hspace{.33em}\\~\\ c.5\hspace{.33em}\\~\\ d.6\hspace{.33em}\\~\\ P,Q,R,S \in \mathbb{Z} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (kainui):

Are we given any other info? Are these all integers or are we given r and s?

OpenStudy (mathmath333):

r and s are integers, no other info is given

OpenStudy (mathmath333):

and p and q also are integers

OpenStudy (kainui):

I am just sort of playing around with it right now, but we can safely say that (P+2) and (Q+2) are either both positive or both negative, which might come in useful.

OpenStudy (mathmath333):

yes i get they can be both \(\pm\) , i think the option are used to solve this

OpenStudy (anonymous):

what is R/s ?

OpenStudy (kainui):

Right, so P and Q are both positive, R and S are also then both positive, which makes it simpler to think about.

OpenStudy (anonymous):

from second ratio we got 2P-Q=-2

OpenStudy (anonymous):

or Q=2(p+1)

mathslover (mathslover):

What about using Componendo and Dividendo?

OpenStudy (mathmath333):

yes thats very useful, lol

OpenStudy (anonymous):

yes @mathslover it works

OpenStudy (anonymous):

so to continue on what im working R/S=1/2 :)

OpenStudy (mathmath333):

@Marki how u got \(\dfrac{r}{s}=1/2\)

OpenStudy (anonymous):

from second ratio Q=2(p+1) so p+1/2(p+1)=R/S=1/2

OpenStudy (michele_laino):

from the proportion: \[(p+2):(q+2)=1:2\] I can write this one: \[(p+2+q+2):(p+2)=-1:1\] from which I get: \[p+q=3p+2\] which is compatible only with option c, sin p has to be an integer number

OpenStudy (michele_laino):

oops..since p has to be....

OpenStudy (michele_laino):

oops..error not -1:1 it is 3:1

OpenStudy (mathmath333):

@Michele_Laino (p+2+q+2):(p+2)=3:1 ?

OpenStudy (michele_laino):

yes

mathslover (mathslover):

Cool. \(\cfrac{P+2}{Q+2} = \cfrac{1}{2} \) Using Componendo and Dividendo. \(\cfrac{P + Q + Q + 2 }{P + 2 - Q - 2 } = \cfrac{1 + 2 }{ 1 - 2 } \) \(\cfrac{P + 2Q + 2}{P - Q} = -3 \)

mathslover (mathslover):

Oops . Mistake : \(\cfrac{P + Q + 2 + 2 }{ P - Q } = -3 \)

OpenStudy (mathmath333):

how do i proceed

mathslover (mathslover):

\( P + Q + 4 = -3(P-Q) \) \(P + Q + 4 = -3 P + 3Q \\ -4P + 2Q = 4 \\ -2P + Q = 2\) \(Q = 2 + 2P \) We need to find P + Q. So, P + Q = 2 + 2P + P = 2 + 3P \(\cfrac{P + 1}{Q + 1} = \cfrac{R}{S} \\ \cfrac{P + 1 + Q + 1 }{ P + 1 - Q - 1 } = \cfrac{R + S}{R - S} \\ \cfrac{P + Q + 2 }{ P - Q } = \cfrac{R+S}{R-S} \) \(\cfrac{2 + 2P + P + 2}{P - 2 - 2P} = \cfrac{R+S}{R-S}\) \(\cfrac{3P + 4}{-P-2} = \cfrac{R+S}{R-S}\) Hmm, still thinking.

OpenStudy (mathmath333):

i think the componenedo dividendo is tedious , it will be wise to use options

mathslover (mathslover):

\( \cfrac{P + 1}{Q + 1} = \cfrac{R}{S} \\ \cfrac{P + 1 + Q + 1}{P + 1 - Q - 1} = \cfrac{R + S}{R - S} \\ \cfrac{P + Q}{P - Q} + \cfrac{2}{P-Q} = \cfrac{R+S}{R-S} \\ \) \( \cfrac{P+Q}{P-Q} = \cfrac{R+S}{R-S} - \cfrac{2}{P-Q} \) -------- (1) And \(\cfrac{P+2}{Q+2} = \cfrac{1}{2} \\ \cfrac{P + Q + 4}{ P -Q } = -3 \\ \) \( \cfrac{P+Q}{P-Q} = -3 - \cfrac{4}{P-Q} \) --------------(2) Equate (1) and (2)

mathslover (mathslover):

See if you get something? Not sure how will we use options though.

mathslover (mathslover):

Let P - Q = x \(\cfrac{R+S}{R-S} - \cfrac{2}{P -Q} = -3 - \cfrac{4}{P-Q} \) \(\cfrac{R+S}{R-S} - \cfrac{2}{x} = -3 - \cfrac{4}{x} \) \(\cfrac{R+S}{R-S} + 3 = \cfrac{2}{x} - \cfrac{4}{x} \) \(\cfrac{R+S}{R-S} + 3 = \cfrac{-2}{x} \) \(x \left( \cfrac{R+S}{R-S} + 3 \right) = -2 \) \(x = \cfrac{-2}{ \cfrac{R+S}{ R-S} + 3} \)

OpenStudy (kainui):

If you do this problem in the future, it is highly recommended to do a substitution A=P+1 and B=Q+1 so that your equations look nicer: \[\Large \frac{A}{B}= \frac{R}{S} \\ \Large \frac{A+1}{B+1}=\frac{1}{2} \\ \Large solve: \ A+B-2\]

OpenStudy (mathmath333):

@Kainui still confused

OpenStudy (kainui):

I'm just saying do this instead so that your algebra looks cleaner, I didn't solve anything.

mathslover (mathslover):

Okay, I seem to get something now.

OpenStudy (mathmath333):

by the way i was thinking of solving \(2p-q=-2\\ p+q=3,4,5,6\) any of those which gives integer is solution https://www.desmos.com/calculator/owi578rpi7

OpenStudy (mathmath333):

i mean to save time

OpenStudy (mathmath333):

still any others methods are appriciated

OpenStudy (kainui):

Wow, the first equation is useless, all it tells us is that it's a ratio of two numbers since we're never given R or S. Throw it away! Make the substitution A=P+2 and B=Q+2 and then plug them in to the second equation to get 2A=B. Remember we are trying to solve P+Q, so we plug these in and get A+B-4 that we need to solve. Use the equation I just found (2A=B) and plug it in to get A+2A-4 then we have 3A-4. So from here we just plug in values for A since the answers are all really small, A=2 gives the smallest positive number which is 3*2-4=2. Not an answer, so try A=3 and we have 3*3-4=5. That's answer C. =)

OpenStudy (mathmath333):

lol i like it @Kainui

OpenStudy (kainui):

I will give medal + fan to someone who can prove to me that the first equation is not completely useless.

OpenStudy (mathmate):

We know that there may be many solutions in integer, but if we are limited to the four choices, we can figure it as follows: \(\dfrac{P+2}{Q+2}=\dfrac{1}{2}=\dfrac{2}{4}=\dfrac{3}{6}=...\), then \(\dfrac{P}{Q}=\{\dfrac{-1}{0}or=\dfrac{0}{2} or=\dfrac{1}{4}=...\}\) which means P+Q=5

OpenStudy (mathmath333):

yes that kind of thing i was looking for

OpenStudy (mathmate):

@Kainui I am still intrigued by the statement: " the first equation is not completely useless." I have the impression that if P,Q,R,S \(\in\) Z, then the second equation implies the first, so the first is superfluous... I wonder...

OpenStudy (kainui):

Yeah, it's sort of weird to have that first equation there, I don't see how that is at all useful. The only reason that it might be useful is that it means P is not equal to -1 maybe? But we already know from the second equation that P+2 and Q+2 are both positive from looking at the answer choices.

OpenStudy (anonymous):

:)

OpenStudy (mathmate):

@Kainui Thx! :)

ganeshie8 (ganeshie8):

\[P+Q \equiv 2 \pmod {3}\]

OpenStudy (anonymous):

Q=2(P+1) P+Q=3P+2 :O

ganeshie8 (ganeshie8):

thats the general solution for given equations

ganeshie8 (ganeshie8):

*diophantine equations

OpenStudy (anonymous):

yeah so least one is 5 >.<

ganeshie8 (ganeshie8):

Oh yes it has to be P+Q = 3P+2